# Robust Multi-Modal Density Estimation

### v1

@@ Abstract
Development of multi-modal, probabilistic prediction models has lead to a need for comprehensive evaluation metrics. While several metrics can characterize the accuracy of machine-learned models (e.g., negative log-likelihood, Jensen-Shannon divergence), these metrics typically operate on probability densities. Applying them to purely sample-based prediction models thus requires that the underlying density function is estimated. However, common methods such as kernel density estimation (KDE) have been demonstrated to lack robustness, while more complex methods have not been evaluated in multi-modal estimation problems. In this paper, we present ROME (RObust Multi-modal density Estimator), a non-parametric approach for density estimation which addresses the challenge of estimating multi-modal, non-normal, and highly correlated distributions. ROME utilizes clustering to segment a multi-modal set of samples into multiple uni-modal ones and then combines simple KDE estimates obtained for individual clusters in a single multi-modal estimate. We compared our approach to state-of-the-art methods for density estimation as well as ablations of ROME, showing that it not only outperforms established methods but is also more robust to a variety of distributions. Our results demonstrate that ROME can overcome the issues of over-fitting and over-smoothing exhibited by other estimators, promising a more robust evaluation of probabilistic machine learning models.
@@ 1 Introduction
The criticality of many modern AI models necessitates a thorough examination of their accuracy. Because of this, arguably any data-driven model can only be taken seriously if demonstrated to sufficiently match the ground truth data. Nevertheless, a significant challenge arises when evaluating probabilistic models that incorporate randomness into their outputs. Examples of such models include generative adversarial networks Goodfellow et al. (2020), variational autoencoders Wei et al. (2020), normalizing flows Kobyzev et al. (2020), and their numerous variations.
@@
When probabilistic models are trained on multi-modal data, they are often evaluated using simplistic metrics (e.g., Mean Squared Error (MSE) from the ground truth). However, such simplistic metrics are unsuited for determining how well a predicted distribution corresponds to the underlying distribution, as they do not capture the fit to the whole distribution. For example, the lowest MSE value could be achieved by accurate predictions of the mean of the true underlying distribution whereas potential differences in variance and shape of the distribution would not be penalized. This necessitates more advanced metrics that evaluate the match between the model and the (potentially multi-modal) data. For instance, Negative Log-Likelihood (NLL), Jensen-Shannon Divergence (JSD), and Expected Calibration Error (ECE) can be used to evaluate how well the full distribution of the data is captured by the learned models Xu et al. (2019); Mozaffari et al. (2020); Rasouli (2020). However, most data-driven models represent the learned distribution implicitly, only providing individual samples and not the full distribution as an output. This complicates the comparison of the model output to the ground-truth data distributions since the above metrics require distributions, not samples, as an input. Practically, this issue can be addressed by estimating the predicted probability density based on samples generated by the model.
Simple methods like Gaussian Mixture Models (GMM), Kernel Density Estimation (KDE) Deisenroth et al. (2020), and k-Nearest Neighbors (kNN) Loftsgaarden and Quesenberry (1965) are commonly used for estimating probability density functions. These estimators however rely on strong assumptions about the underlying distribution, and can thereby introduce biases or inefficiencies in the estimation process. For example, problems can arise when encountering multi-modal, non-normal, and highly correlated distributions (see Section 2). While more advanced methods such as Manifold Parzen Windows (MPW) Vincent and Bengio (2002) and Vine Copulas (VC) Nagler and Czado (2016) exist, they have not been thoroughly tested on such problems, which questions their performance.
@@ 2 Related Work
The most common class of density estimators are so-called Parzen windows Parzen (1962), which estimate the density through a combination of parametric Probability Density Functions (PDFs). A number of common methods use this approach, with KDE being a common non-parametric method Silverman (1998). Provided a type of kernel – which is oftentimes a Gaussian but can be any other type of distribution – KDE places the kernels around each sample of a data distribution and then sums over these kernels to get the final density estimation over the data. This method is often chosen as it does not assume any underlying distribution type Silverman (1998). However, if the underlying distribution is highly correlated, then the common use of a single isotropic kernel function can lead to over-smoothing Vincent and Bengio (2002); Wang et al. (2009). Among multiple approaches for overcoming this issue Wang et al. (2009); Gao et al. (2022), especially noteworthy is the MPW approach Vincent and Bengio (2002). It uses a unique anisotropic kernel for every datapoint, estimated based on the correlation of the direct neighbors of each sample. However, it has not been previously tested in high-dimensional benchmarks, which is especially problematic as the required memory scales quadratically with the dimensionality of the problem.
@@ 3 RObust Multi-modal Estimator (ROME)
The problem of density estimation can be formalized as finding queryable $\widehat{p}\in\mathcal{P}$, where
@@
such that $\widehat{p}$ is close in some sense to the non-queryable PDF $p$ underlying the $N$ available $M$-dimensional samples $\bm{X}\in\mathbb{R}^{N\times M}$: $\bm{X}\sim p$.
A solution to the above problem would be an estimator $f:\mathbb{R}^{N\times M}\rightarrow\mathcal{P}$, resulting in $\widehat{p}=f(\bm{X})$. Our proposed estimator $f_{\text{ROME}}$ (Algorithm 1) is built on top of non-parametric cluster extraction. Namely, by separating groups of samples surrounded by areas of low density – expressing the mode of the underlying distribution – we reduce the multi-modal density estimation problem to multiple uni-modal density estimation problems for each cluster. The distributions within each cluster then become less varied in density or correlation than the full distribution. Combining this with decorrelation and normalization, the use of established methods such as KDE to estimate probability densities for those uni-modal distributions is now more promising, as problems with multi-modality and correlated modes (see Section 2) are accounted for. The complete multi-modal distribution is then obtained as a weighted average of the estimated uni-modal distributions.
@@ 3.1 Extracting Clusters
To cluster samples $\bm{X}$, ROME employs the OPTICS algorithm Ankerst et al. (1999) that can detect clusters of any shape with varying density. This algorithm uses reachability analysis to sequentially transfer samples from a set of unincluded samples $\bm{X}_{U,i}$ to the set of included samples $\bm{X}_{I,i}$, starting with a random sample $\bm{x}_{1}$ ($\bm{X}_{I,1}=\{\bm{x}_{1}\}$ and $\bm{X}_{U,1}=\bm{X}\backslash\{\bm{x}_{1}\}$). The samples $\bm{x}_{i}$ are then selected at iteration $i$ in the following way:
$\bm{x_{i}}=\underset{\bm{x}\in\bm{X}_{U,i-1}}{\operatorname*{arg\,min}}\,r_{i}(\bm{x})=\underset{\bm{x}\in\bm{X}_{U,i-1}}{\operatorname*{arg\,min}}\;\,\underset{\bm{\widetilde{x}}\in\bm{X}_{I,i-1}}{\min\vphantom{g}}\,d_{r}\left(\bm{x},\bm{\widetilde{x}}\right)\,.$ (1)
This sample is then transferred between sets, with $\bm{X}_{I,i}=\bm{X}_{I,i-1}\cup\{\bm{x}_{i}\}$ and $\bm{X}_{U,i}=\bm{X}_{U,i-1}\backslash\{\bm{x}_{i}\}$, while expanding the reachability set $\bm{R}_{i}=\bm{R}_{i-1}\cup\left\{r_{i}(\bm{x}_{i})\right\}$ (with $\bm{R}_{1}=\left\{\infty\right\}$). The reachability distance $d_{r}$ in Equation (1) is defined as
@@
where $k_{\bm{c}}$ is:
Here, $k_{\min}$ is needed to ensure that the method is stable, as a too low $k_{\bm{c}}$ would make the subsequent clustering vulnerable to sampling fluctuations. Meanwhile, $k_{\max}$ ensures that the reachability distances are actually based on only local information, and are not including points from other modes. Lastly, the term $NM/\alpha_{k}$ is used to keep $r_{i}$ (see Equation (1)) independent from the number of samples, while allowing for the higher number of samples needed in higher-dimensional spaces.
The next part of the OPTICS algorithm – after obtaining the reachability distances $\bm{R}_{N}$ and the ordered set $\bm{X}_{I,N}$ – is the extraction of a set of clusters $\bm{C}$, with clusters $C=\{c_{\min},\ldots,c_{\max}\}\in\bm{C}$ (with $\bm{X}_{C}=\{\bm{x}_{I,N,j}\,|\,j\in C\}\in\mathbb{R}^{|C|\times M}$ and $|C|\geq 2$). As the computational cost of creating such a cluster set is negligible compared to the reachability analysis, we can test multiple clusterings generated using two different approaches (with $r_{\text{bound}}=\min\{r_{N,c_{\min}},r_{N,c_{\max}+1}\}$):
  • First, we use DBSCAN Ester et al. (1996) for clustering based on an absolute limit $\epsilon$, where a cluster must fulfill the condition:
@@
  • Second, we use $\bm{\xi}$-clustering Ankerst et al. (1999) to cluster based on a proportional limit $\xi$, where a cluster fulfills:
@@
Samples not included in any cluster $C$ that fulfills the respective conditions above are kept in a separate noise cluster $C_{\text{noise}}\in\bm{C}$. Upon generating different sets of clusters $\bm{C}$ for $\epsilon\in\bm{\varepsilon}$ and $\xi\in\bm{\xi}$, we select the $\bm{C}$ that achieves the highest silhouette score [^3] The silhouette score measures the similarity of each object to its own cluster’s objects compared to the other clusters’ objects. $S=\text{{SIL}}(\bm{C},\bm{X}_{I,N})\in[-1,1]$ Rousseeuw (1987). The clustering then allows us to use PDF estimation methods on uni-modal distributions.
@@ 3.3 Normalization
Here, $\widetilde{\bm{\Sigma}}_{C}$ is a diagonal matrix with the entries $\widetilde{\sigma}_{C}^{(m)}$. To avoid over-fitting to highly correlated distributions, we introduce a regularization with a value $\sigma_{\min}$ (similar to Vincent and Bengio (2002)) that is applied to the empirical standard deviation $\sigma_{\text{PCA},C}^{(m)}=\sqrt{\mathbb{V}_{m}\left(\bm{X}_{\text{PCA},C}\right)}$ ($\mathbb{V}_{m}$ is the variance along feature $m$) of each rotated feature $m$:
@@ 3.4 Estimating the Probability Density Function
The bandwidth $b_{C}$ is then set using Silverman’s rule Silverman (1998):
@@
Here, the term $|C|/N$ is used to weigh the different distributions of each cluster with the size of each cluster, so that each sample is represented equally. As the different KDEs $\widehat{p}_{C}$ are fitted to the transformed samples, we apply them not to the original sample $\bm{x}$, but instead apply the equal transformation used to generate those transformed samples, using $\widehat{p}_{C}\left(\left(\bm{x}-\overline{\bm{x}}_{C}\right)\right)$. To account for the change in density within a cluster $C$ introduced by this transformation, we use the factor $|\det\left(\bm{T}_{C}\right)|$.
@@
Distribution
ROME MPW VC ROME MPW VC ROME MPW VC
Aniso
$0.010\scriptscriptstyle{\pm 0.001}$
$0.026\scriptscriptstyle{\pm 0.002}$
$\underline{0.005}\scriptscriptstyle{\pm 0.001}$
$\underline{-0.13}\scriptscriptstyle{\pm 0.31}$
$-0.60\scriptscriptstyle{\pm 0.13}$
$1.91\scriptscriptstyle{\pm 0.91}$
$\underline{-2.53}\scriptscriptstyle{\pm 0.02}$
$-2.57\scriptscriptstyle{\pm 0.02}$
$-3.19\scriptscriptstyle{\pm 0.02}$
Varied
$0.011\scriptscriptstyle{\pm 0.001}$
$0.025\scriptscriptstyle{\pm 0.002}$
$\underline{0.008}\scriptscriptstyle{\pm 0.001}$
$\underline{-0.13}\scriptscriptstyle{\pm 0.20}$
$-0.49\scriptscriptstyle{\pm 0.11}$
$1.27\scriptscriptstyle{\pm 0.53}$
$\underline{-4.10}\scriptscriptstyle{\pm 0.03}$
$-4.12\scriptscriptstyle{\pm 0.03}$
$-4.29\scriptscriptstyle{\pm 0.03}$
Two Moons
$\underline{0.002}\scriptscriptstyle{\pm 0.001}$
$0.023\scriptscriptstyle{\pm 0.002}$
$0.008\scriptscriptstyle{\pm 0.002}$
$1.40\scriptscriptstyle{\pm 0.52}$
$\underline{-0.52}\scriptscriptstyle{\pm 0.10}$
$1.36\scriptscriptstyle{\pm 0.51}$
$-1.02\scriptscriptstyle{\pm 0.01}$
$\underline{-0.36}\scriptscriptstyle{\pm 0.01}$
$-0.95\scriptscriptstyle{\pm 0.01}$
Trajectories
$\underline{0.008}\scriptscriptstyle{\pm 0.002}$
$0.016\scriptscriptstyle{\pm 0.001}$
$0.743\scriptscriptstyle{\pm 0.005}$
$\underline{1.03}\scriptscriptstyle{\pm 0.22}$
$1.24\scriptscriptstyle{\pm 0.24}$
$9.30\scriptscriptstyle{\pm 1.30}$
$\underline{29.32}\scriptscriptstyle{\pm 0.02}$
$26.09\scriptscriptstyle{\pm 0.02}$
$-215.23\scriptscriptstyle{\pm 17.6}$

### v2

# ROME: Robust Multi-Modal Density Estimator [^1] Extended version with Appendix: https://arxiv.org/abs/2401.10566

@@ Abstract
The estimation of probability density functions is a fundamental problem in science and engineering. However, common methods such as kernel density estimation (KDE) have been demonstrated to lack robustness, while more complex methods have not been evaluated in multi-modal estimation problems. In this paper, we present ROME (RObust Multi-modal Estimator) [^3] Source code: https://github.com/anna-meszaros/ROME, a non-parametric approach for density estimation which addresses the challenge of estimating multi-modal, non-normal, and highly correlated distributions. ROME utilizes clustering to segment a multi-modal set of samples into multiple uni-modal ones and then combines simple KDE estimates obtained for individual clusters in a single multi-modal estimate. We compared our approach to state-of-the-art methods for density estimation as well as ablations of ROME, showing that it not only outperforms established methods but is also more robust to a variety of distributions. Our results demonstrate that ROME can overcome the issues of over-fitting and over-smoothing exhibited by other estimators.
@@ 1 Introduction
Numerous processes are non-deterministic by nature, from geological and meteorological occurrences, biological activities, as well as the behavior of living beings. Estimating the underlying probability density functions (PDFs) of such processes enables a better understanding of them and opens possibilities for probabilistic inference regarding future developments. Density estimation is instrumental in many applications including classification, anomaly detection, and evaluating probabilistic AI models, such as generative adversarial networks Goodfellow et al. (2020), variational autoencoders Wei et al. (2020), normalizing flows Kobyzev et al. (2020), and their numerous variations.
@@
When probabilistic models are trained on multi-modal data, they are often evaluated using simplistic metrics (e.g., mean squared error (MSE) between the predicted and ground truth samples). However, such simplistic metrics are unsuited for determining how well a predicted distribution corresponds to the underlying distribution, as they do not capture the fit to the whole distribution. For example, the lowest MSE value between true and predicted samples could be achieved by accurate predictions of the mean of the true underlying distribution whereas potential differences in variance or shape of the distribution would not be penalized. This necessitates more advanced metrics that evaluate the match between the model and the (potentially multi-modal) data. For instance, negative log-likelihood (NLL), Jensen-Shannon divergence (JSD), and expected calibration error (ECE) can be used to evaluate how well the full distribution of the data is captured by the learned models Xu et al. (2019); Mozaffari et al. (2020); Rasouli (2020). However, most data-driven models represent the learned distribution implicitly, only providing individual samples and not the full distribution as an output. This complicates the comparison of the model output to the ground-truth data distributions since the above metrics require distributions, not samples, as an input. Practically, this can be addressed by estimating the predicted probability density based on samples generated by the model.
Simple methods like Gaussian mixture models (GMM), kernel density estimation (KDE) Deisenroth et al. (2020), and k-nearest neighbors (kNN) Loftsgaarden and Quesenberry (1965) are commonly used for estimating probability density functions. These estimators however rely on strong assumptions about the underlying distribution, and can thereby introduce biases or inefficiencies in the estimation process. For example, problems can arise when encountering multi-modal, non-normal, and highly correlated distributions (see Section 2). While more advanced methods such as manifold Parzen windows (MPW) Vincent and Bengio (2002) and vine copulas (VC) Nagler and Czado (2016) exist, they have not been thoroughly tested on such problems, which raises questions about their performance.
@@ 2 Related Work
The most common class of density estimators are so-called Parzen windows Parzen (1962), which estimate the density through an aggregation of parametric probability density functions. A number of common methods use this approach, with KDE being a common non-parametric method Silverman (1998). Provided a type of kernel – which is oftentimes a Gaussian but can be any other type of distribution – KDE places the kernels around each sample of a data distribution and then sums over these kernels to get the final density estimation over the data. This method is often chosen as it does not assume any underlying distribution type Silverman (1998). However, if the underlying distribution is highly correlated, then the common use of a single isotropic kernel function can lead to over-smoothing Vincent and Bengio (2002); Wang et al. (2009). Among multiple approaches for overcoming this issue Wang et al. (2009); Gao et al. (2022), especially noteworthy is the MPW approach Vincent and Bengio (2002). It uses a unique anisotropic kernel for every datapoint, estimated based on the correlation of the k-nearest neighbors of each sample. However, it has not been previously tested in high-dimensional benchmarks, which is especially problematic as the required memory scales quadratically with the dimensionality of the problem.
@@ 3 RObust Multi-modal Estimator (ROME)
The problem of density estimation can be formalized as finding a queryable $\widehat{p}\in\mathcal{P}$, where
@@
such that $\widehat{p}$ is close to the non-queryable PDF $p$ underlying the $N$ available $M$-dimensional samples $\bm{X}\in\mathbb{R}^{N\times M}$: $\bm{X}\sim p$.
A solution to the above problem would be an estimator $f:\mathbb{R}^{N\times M}\rightarrow\mathcal{P}$, resulting in $\widehat{p}=f(\bm{X})$. Our proposed estimator $f_{\text{ROME}}$ (Algorithm 1) is built on top of non-parametric cluster extraction. Namely, by separating groups of samples surrounded by areas of low density – expressing the mode of the underlying distribution – we reduce the multi-modal density estimation problem to multiple uni-modal density estimation problems for each cluster. The distributions within each cluster then become less varied in density or correlation than the full distribution. Combining this with decorrelation and normalization, the use of established methods such as KDE to estimate probability densities for those uni-modal distributions is now more promising, as problems with multi-modality and correlated modes (see Section 2) are accounted for. The multi-modal distribution is then obtained as a weighted average of the estimated uni-modal distributions.
@@ 3.1 Extracting Clusters
To cluster samples $\bm{X}$, ROME employs the OPTICS algorithm Ankerst et al. (1999) that can detect clusters of any shape with varying density using a combination of reachability analysis – which orders the data in accordance to reachability distances – followed by clustering the ordered data based on these reachability distances.
In the first part of the algorithm, the reachability analysis is used to sequentially transfer samples from a set of unincluded samples $\bm{X}_{U,i}$ to the set of included and ordered samples $\bm{X}_{I,i}$, starting with a random sample $\bm{x}_{1}$ ($\bm{X}_{I,1}=\{\bm{x}_{1}\}$ and $\bm{X}_{U,1}=\bm{X}\backslash\{\bm{x}_{1}\}$). The samples $\bm{x}_{i+1}$ are then selected at iteration $i$ based on the reachability distance $r$:
$\bm{x}_{i+1}=\underset{\bm{x}\in\bm{X}_{U,i}}{\operatorname*{arg\,min}}\,r(\bm{x},\bm{X}_{I,i})=\underset{\bm{x}\in\bm{X}_{U,i}}{\operatorname*{arg\,min}}\;\,\underset{\bm{\widetilde{x}}\in\bm{X}_{I,i}}{\min\vphantom{g}}\,d_{r}\left(\bm{x},\bm{\widetilde{x}}\right)\,.$ (1)
This sample is then transferred between sets, with $\bm{X}_{I,i+1}=\bm{X}_{I,i}\cup\{\bm{x}_{i+1}\}$ and $\bm{X}_{U,{i+1}}=\bm{X}_{U,i}\backslash\{\bm{x}_{i+1}\}$, while expanding the reachability set $\bm{R}_{i+1}=\bm{R}_{i}\cup\left\{r(\bm{x}_{i+1},\bm{X}_{I,i})\right\}$ (with $\bm{R}_{1}=\left\{\infty\right\}$). The reachability distance $d_{r}$ in Equation (1) is defined as
@@
where ${\min}_{k_{\bm{c}}}$ is the $k_{\bm{c}}$-smallest value from all the available $\bm{\widehat{x}}$, which is used to smooth out random local density fluctuations. We use
$k_{\bm{c}}=\min\left\{k_{\max},\max\left\{k_{\min},\frac{NM}{\alpha_{k}}\right\}\right\}\,,$ (2)
where $k_{\min}$ and $k_{\max}$ ensure that there are always sufficient but not too many points used for this smoothing. Meanwhile, the term $NM/\alpha_{k}$ adjusts $k_{\bm{c}}$ to the number of samples and dimensions.
The second part of the OPTICS algorithm – after obtaining the reachability distances $\bm{R}_{N}$ and the ordered set $\bm{X}_{I,N}$ – is the extraction of a set of clusters $\bm{C}$, with clusters $C=\{c_{\min},\ldots,c_{\max}\}\in\bm{C}$ (with $\bm{X}_{C}=\{\bm{x}_{I,N,j}\,|\,j\in C\}\in\mathbb{R}^{|C|\times M}$). As the computational cost of creating such a cluster set is negligible compared to the reachability analysis, we can test multiple clusterings generated using two different approaches (with $r_{\text{bound}}=\min\{r_{N,c_{\min}},r_{N,c_{\max}+1}\}$, see Appendix A for further discussion):
  • First, we use DBSCAN Ester et al. (1996) for generating the clustering $\bm{C}_{\epsilon}$ based on an absolute limit $\epsilon$, where a cluster must fulfill the condition:
@@
  • Second, we use $\bm{\xi}$-clustering Ankerst et al. (1999) to generate the clustering $\bm{C}_{\xi}$ based on the proportional limit $\xi$, where a cluster fulfills:
@@
In both cases, each prospective cluster $C$ also has to fulfill the condition $|C|\geq 2$. However, it is possible that not every sample can be fit into a cluster fulfilling the conditions above. These samples are then kept in a separate noise cluster $C_{\text{noise}}$ that does not have to fulfill those conditions ($C_{\text{noise}}\in\bm{C}_{\epsilon}$ or $C_{\text{noise}}\in\bm{C}_{\xi}$ respectively). Upon generating multiple sets of clusters $\bm{C}_{\epsilon}$ ($\epsilon\in\bm{\varepsilon}$) and $\bm{C}_{\xi}$ ($\xi\in\bm{\xi}$), we select the final set of clusters $\bm{C}$ that achieves the highest silhouette score [^4] The silhouette score measures the similarity of each object to its own cluster’s objects compared to the other clusters’ objects. $S=\text{{SIL}}(\bm{C},\bm{X}_{I,N})\in[-1,1]$ Rousseeuw (1987). The clustering then allows us to use PDF estimation methods on uni-modal distributions.
@@ 3.3 Normalization
Here, $\widetilde{\bm{\Sigma}}_{C}$ is a diagonal matrix with the entries $\widetilde{\sigma}_{C}^{(m)}$. To avoid over-fitting to highly correlated distributions, we introduce a regularization with a value $\sigma_{\min}$ (similar to Vincent and Bengio (2002)) that is applied to the empirical standard deviation $\sigma_{\text{PCA},C}^{(m)}=\sqrt{\mathbb{V}_{m}\left(\bm{X}_{\text{PCA},C}\right)}$ ($\mathbb{V}_{m}$ is the variance along feature $m$) of each rotated feature $m$ (with $\bm{C}^{\prime}=\bm{C}\backslash\{C_{\text{noise}}\}$):
@@ 3.4 Estimating the Probability Density Function
The bandwidth $b_{C}$ is set using Silverman’s rule Silverman (1998):
@@
Here, the term $|C|/N$ is used to weigh the different distributions of each cluster with the size of each cluster, so that each sample is represented equally. As the different KDEs $\widehat{p}_{C}$ are fitted to the transformed samples, we apply them not to the original sample $\bm{x}$, but instead apply the identical transformation used to generate those transformed samples, using $\widehat{p}_{C}\left(\left(\bm{x}-\overline{\bm{x}}_{C}\right)\right)$. To account for the change in density within a cluster $C$ introduced by the transformation $\bm{T}_{C}$, we use the factor $|\det\left(\bm{T}_{C}\right)|$.
@@
Distribution ROME MPW VC ROME MPW VC ROME MPW VC
Aniso $0.010\scriptscriptstyle{\pm 0.001}$ $0.026\scriptscriptstyle{\pm 0.002}$ $\underline{0.005}\scriptscriptstyle{\pm 0.001}$ $\underline{-0.13}\scriptscriptstyle{\pm 0.31}$ $-0.60\scriptscriptstyle{\pm 0.13}$ $1.91\scriptscriptstyle{\pm 0.91}$ $\underline{-2.53}\scriptscriptstyle{\pm 0.02}$ $-2.57\scriptscriptstyle{\pm 0.02}$ $-3.19\scriptscriptstyle{\pm 0.02}$
Varied $0.011\scriptscriptstyle{\pm 0.001}$ $0.025\scriptscriptstyle{\pm 0.002}$ $\underline{0.008}\scriptscriptstyle{\pm 0.001}$ $\underline{-0.13}\scriptscriptstyle{\pm 0.20}$ $-0.49\scriptscriptstyle{\pm 0.11}$ $1.27\scriptscriptstyle{\pm 0.53}$ $\underline{-4.10}\scriptscriptstyle{\pm 0.03}$ $-4.12\scriptscriptstyle{\pm 0.03}$ $-4.29\scriptscriptstyle{\pm 0.03}$
Two Moons $\underline{0.002}\scriptscriptstyle{\pm 0.001}$ $0.023\scriptscriptstyle{\pm 0.002}$ $0.008\scriptscriptstyle{\pm 0.002}$ $1.40\scriptscriptstyle{\pm 0.52}$ $\underline{-0.52}\scriptscriptstyle{\pm 0.10}$ $1.36\scriptscriptstyle{\pm 0.51}$ $-1.02\scriptscriptstyle{\pm 0.01}$ $\underline{-0.36}\scriptscriptstyle{\pm 0.01}$ $-0.95\scriptscriptstyle{\pm 0.01}$
Trajectories $\underline{0.008}\scriptscriptstyle{\pm 0.002}$ $0.016\scriptscriptstyle{\pm 0.001}$ $0.743\scriptscriptstyle{\pm 0.005}$ $\underline{1.03}\scriptscriptstyle{\pm 0.22}$ $1.24\scriptscriptstyle{\pm 0.24}$ $9.30\scriptscriptstyle{\pm 1.30}$ $\underline{29.32}\scriptscriptstyle{\pm 0.02}$ $26.09\scriptscriptstyle{\pm 0.02}$ $-215.23\scriptscriptstyle{\pm 17.6}$